# Addressing Source Scale Bias via Image Warping for Domain Adaptation

### v1

@@ Appendix F Additional Technical Details
Ground Truth Segmentation to Boxes: Instance-based image warping requires bounding boxes, which are not provided in some semantic segmentation benchmarks like GTA [28]. To circumvent this, we generate “from-seg” bounding boxes from ground truth semantic segmentation maps. We do this by identifying connected components that represent individual instances of foreground categories — specifically, traffic lights, traffic signs, persons, riders, cars, trucks, buses, trains, motorcycles, and bikes.
For each “from-seg” instance, we compute the bounding boxes by finding the minimum enclosing axis-aligned rectangle. These “from-seg” bounding boxes are then employed for instance-level image warping in a similar manner to ground truth bounding boxes.

### v2

@@ Appendix G Additional Technical Details
Ground Truth Segmentation to Boxes: Instance-level saliency guided image warping requires bounding boxes, which are not provided in some semantic segmentation benchmarks like GTA [26]. To address this, we generate ‘from-seg’ bounding boxes from ground truth semantic segmentation maps. Specifically, we first identify connected components representing individual instances of foreground categories, including traffic lights, traffic signs, persons, riders, cars, trucks, buses, trains, motorcycles, and bikes. For each ‘from-seg’ instance, we then compute the bounding boxes by finding the minimum enclosing axis-aligned rectangle. These ‘from-seg’ bounding boxes are finally used for instance-level saliency guidance in the same way as ground truth bounding boxes.